2. A car on a hydraulic lift rests on a cylinder with a piston of a diameter of 12 inches. If a connecting cylinder with a piston of i -inch diameter is driven by compressed air with a force of 100 lb., how big the force approximately will be generated to the larger piston to lift the car?

a. 600 lb. b.1,200 lb. c. 6,000 lb. d. 12,000 lb. e. 15,000 lb


thats all the informtion that was given to this problem
Answer #1

Force is given by:

Force = Pressure*Area

Pressure = Force/Area

Pressure on both piston will be equal, So

F1/A1 = F2/A2

F2 = F1*(A2/A1) = F1*(d2/d1)^2

Area of piston = pi*d^2/4

where d = diameter of piston

F1 = Force by which smaller piston is pressed = 100 lb

d2 = diameter of larger piston = 12 inches

d1 = diameter of smaller piston = 1 inches (Please check again your question, it should be 1 inch, if there is some other value let me know through comments)

So,

F2 = (100 lb)*(12/1)^2

F2 = 100*144 = 14400 lb

F2 = Approximately 15000 lb

Correct option is E.

Please Upvote.